1. Synthesis of compound in the gaseous phase reaction has been studied at two temperatures. At 298 K, the kp value is 1.22×10-3 and at 498 K the kp value is 2.16. Calculate the standard enthalpy change of this chemical reaction. Where value of R is 1.987Cal K-1mol-1.


1
Expert's answer
2021-05-12T06:27:24-0400

1.987 Cal"\times"K-1"\times"mol-1 = 8.314 J"\times"mol-1"\times"K-1

"\Delta G = -RTlnK_p"

"\Delta G_1 = -RT_1lnK_{p1}=-8.314\times298\times ln(1.22\times10^{-3}) = 16621.8"

"\Delta G_2 = -RT_2lnK_{p2}=-8.314\times498\times ln(2.16) = -3188.5"

"\Delta G = \Delta H - T\Delta S"

"\Delta H = \Delta G + T\Delta S"

"\Delta S = \frac{\Delta H - \Delta G}{T} = \frac{\Delta H - \Delta G_1}{T_1}"

"\Delta H = \Delta G_2 + T_2(\frac{\Delta H - \Delta G_1}{T_1}) = \Delta G_2 + \frac{T_2\Delta H}{T_1} - \frac{T_2\Delta G_1}{T_1}"

"T_1 \Delta H = T_1\Delta G_2 + T_2\Delta H - T_2\Delta G_1"

"T_1 \Delta H -T_2\Delta H = T_1\Delta G_2 - T_2\Delta G_1"

"\Delta H(T_1-T_2) = T_1\Delta G_2 - T_2\Delta G_1"

"\Delta H = \frac{T_1\Delta G_2 - T_2\Delta G_1}{(T_1-T_2)}"

"\Delta H = \frac{298\times(-3188.5)-498\times(16621.8)}{298-498}=46139 \ (J)"


Answer: "\Delta"H = 46139 J.